Subtraction problem
Find the missing values of A, B and C where A>B and B>C in the following subtraction problem. ABC-CBA=CAB
I need help on this quick!

Found 2 solutions by khwang, kolay:
Answer by khwang(438)

ABC
- CBA
-----
= CAB
A > B > C implies
So, 1) A+ B = 10 +C
A + B = 9 + B so A = 9
& A - 1 = C so C = 8 and B = 9
Answer
998
-)899
------
099

Something wrong with the given question, plz double check it.
Kenny

Answer by kolay(1)

954
- 459
-------
495